OpenStudy (anonymous):

how do you solve cosx+1/sin^3=cscx/1-cosx

OpenStudy (psymon):

Kinda tricky, but I'll try and hint you in the right direction and see where we can get. Start by making sin^3(x) into sinx(sin^2(x)). Then change sin^2(x) into cosines using this identity. \[\sin ^{2}x + \cos ^{2}x + 1\] If you are able to correctly replace sin^2(x), then you'll be able to do some factoring in the denominator with difference of squares to get things to cancel.

OpenStudy (anonymous):

ok you lost me wen you wrote that sin^2x changes to what you wrote after

OpenStudy (psymon):

\[(a ^{2}-b ^{2}) = (a-b)(a+b)\]

OpenStudy (anonymous):

how do I use the identity

OpenStudy (psymon):

Well, the identity above lets you do a substitution. Since we know that \[\sin ^{2}x + \cos ^{2}x = 1\], we can solvefor sin^2(x) and then replace it. So when you isolate sin^2(x), what do you have on the other side of the equal sign?

OpenStudy (anonymous):

cos^2x???

OpenStudy (psymon):

1-cos^2(x)

OpenStudy (anonymous):

oh yah

OpenStudy (psymon):

Subtract cos^2(x) from both sides and you see that sin^2(x) is the exact same thing as 1-cos^2(x). Because of that I can rewrite what we have to become: \[\frac{ cosx+1 }{ sinx(1-\cos ^{2}x) }\] Now the cosine part on bottom is a difference of squares. Use the fact that \[(a ^{2}-b ^{2}) = (a-b)(a+b)\] to help you finish.

OpenStudy (anonymous):

so would I hav to use that on the cos^2

OpenStudy (psymon):

Correct.

OpenStudy (anonymous):

what would I hav to do after I have factored cos^2

OpenStudy (psymon):

What'd you factor 1-cos^2(x) into?

OpenStudy (anonymous):

idk

OpenStudy (psymon):

Well 1-cos^2(x) fits the form above of a^2-b^2. So that means you can turn it into (a-b)(a+b). So what is a?

OpenStudy (anonymous):

1???

OpenStudy (psymon):

Yep. So sqrt of 1 is 1, lol. So b would be cosx, right?

OpenStudy (anonymous):

yes

OpenStudy (psymon):

a^2 = 1 b^2 = cos^2(x) a = 1 b = cosx (a^2-b^2) = (a-b)(a+b) therefore (1^2 - cos^2(x)) = (1+cosx)(1-cosx) Kinda see what I did?

OpenStudy (anonymous):

yah I see

OpenStudy (psymon):

Awesome. SO that leaves us with: \[\frac{ cosx+1 }{ (sinx)(cosx+1)(cosx-1) }\]

OpenStudy (anonymous):

yes/ok

OpenStudy (psymon):

cosx + 1 cancels and we have. \[\frac{ 1 }{ sinx }*\frac{ 1 }{ 1-cosx }\] I wrote it backwards last time, my bad. This is the way it should be, lol.

OpenStudy (psymon):

Now just change 1/sin into csc :P

OpenStudy (anonymous):

so wat would happen to the last cosx-1

OpenStudy (anonymous):

never mind I know it would be csx / 1-cosx

OpenStudy (psymon):

Right xD

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

would I be able to ask you one last question???

OpenStudy (psymon):

go for it.

OpenStudy (anonymous):

but its a different problem would that be ok

OpenStudy (psymon):

Yeah, Ill see if I can do it.

OpenStudy (anonymous):

ok well its this 1-tan^2x / 1+tan^2x=1-2sin^2x

OpenStudy (psymon):

This problem uses two identities. One of them is just a transformation of one we used before and the other one is a more obscure one. If you have to do a lot ofthese identity problerms, I suggest having formulas handy.

OpenStudy (anonymous):

I got some

OpenStudy (anonymous):

*formulas

OpenStudy (psymon):

Alright, cool. Well the first identity we're going to use is a transformation of: \[\sin ^{2}x + \cos ^{2}x = 1 \]What I'm going to do is divide everything by cos^2(x). This will just modify the identity into different trig functions, but is still valid: \[\frac{ \sin ^{2}x }{ \cos ^{2}x }+\frac{ \cos ^{2}x }{ \cos ^{2}x }= \frac{ 1 }{ \cos ^{2}x }\]this becomes \[\tan ^{2}x + 1 = \sec ^{2}x\] Now as you should notice, you have something like this result directly in your problem :P

OpenStudy (anonymous):

do you think you could draw it

OpenStudy (psymon):

Yeah, I can. This is just showing where the first identity we use comes from, though.

OpenStudy (psymon):



OpenStudy (anonymous):

ok

OpenStudy (psymon):

mhm. So from there I'm going to split our problem into 2 fractions like this: \[\frac{ 1 }{ \sec ^{2}x }-\frac{ \tan ^{2}x }{ \sec ^{2}x }\] Now from here I can change everything into sines and cosines. Would you know how to do that?

OpenStudy (anonymous):

sec would change to 1/cosx and the tan to sinx/cosx

OpenStudy (anonymous):

right

OpenStudy (anonymous):

is this the problem or is this the way it would sort of be??

OpenStudy (psymon):

This is me going through the problem. Im just showing you step by step. The first step was changing the bottom into sec^2(x), then we separated the fractions. Now we're turning everything into sines and cosines. So if we do what you said, we end up with: \[\cos ^{2}x-\frac{ \sin ^{2}xcos ^{2}x }{ \cos ^{2}x }\] Remember, secx is 1/cosx, but that also means 1/secx is cosx. That is why we have the cos^2(x) in the numerator of both fractions, we had to flip them up. Can you see how that happened?

OpenStudy (anonymous):

umm I cant tell how you set up the equation once you turned everything into sines and cosines

OpenStudy (anonymous):

after you turned 1+tan^x into sec^2x did you keep the top which is 1-tan^2x??

OpenStudy (psymon):

Ah. Well, for every term you have in the numerator, you can make itinto it's own fraction. So I took the term of 1 and made it the numerator of 1 fraction and the -tan^2(x) and made it the numerator of a 2nd fraction. For example: \[\frac{ 1+x }{ 3+y } = \frac{ 1 }{ 3+y }+\frac{ x }{ 3+y }\] This is exactly what I did with our equation. We had: \[\frac{ 1-\tan ^{2}x }{ 1+\tan ^{2}x }\]I turned this into two fractions, one for each term in the numerator: \[\frac{ 1 }{ 1+\tan ^{2}x }- \frac{ \tan ^{2}x }{ 1+\tan ^{2}x }\]Using the identity we opened with, I turned the 1+tan^2(x) into sec^2(x) \[\frac{ 1 }{ \sec ^{2}x }-\frac{ \tan ^{2}x }{ \sec ^{2}x }\] From there, we know that secx is the same as 1/cosx. But since we have 1/secx instead, we say that 1/secx is just cosx. Doing that allows me to rewrite the sec^2(x) parts like so: \[\cos ^{2}x-\tan ^{2}xcos ^{2}x\]Finally, we know tanx is the same as sinx/cosx, so rewriting that I have: \[\cos ^{2}x-\frac{ \sin ^{2}xcos ^{2}x }{ \cos ^{2}x }\]

OpenStudy (anonymous):

I cant tell how you set this up \[\cos ^{2}x-\frac{ \sin ^{2}xcos ^{2}x }{ \cos ^{2}x }\]

OpenStudy (psymon):